OpenStudy (anonymous):

Given a triangle with b = 7, c = 3 , and A = 37° what is the length of a? Round to the nearest tenth. I don't know I can use the law of sines with this, because the length is unavailable for the angle given. I just don't know how I can apply the law of sines to this.

OpenStudy (jtvatsim):

You are right! So, Law of Cosines to the rescue!

OpenStudy (anonymous):

>_< ugh how do I do that?

OpenStudy (merchandize):

Using the law of cosines: a^2 = 7^2 + 3^2 - 2(7)(3)cos37° a^2 = 49 + 9 - 42cos37° a^2 = 58 - 42cos37° a = sqrt (58 - 42cos37°) a = 4.9

OpenStudy (jtvatsim):

\[a^2 = b^2 + c^2 - 2bc\cos{A}\]

OpenStudy (jtvatsim):

It's a mess, but that is the equation to use.

OpenStudy (jtvatsim):

It's like a modified version of the Pythagorean theorem, it works for all triangles though not just right triangles.

OpenStudy (anonymous):

Okai I think I got it c: Thank you.